OpenStudy (anonymous):

I need a doctor for this question~! If 2% of the batteries manufactured by a company are defective, find the probability that in case of 144 batteries, there are 3 defective ones. Tq ^^

OpenStudy (amistre64):

p = .02 q = .98 has something to do with those I think

OpenStudy (anonymous):

is it by using binomial distribution?

OpenStudy (amistre64):

i cant recall :)

OpenStudy (amistre64):

prolly

OpenStudy (jamesj):

Yes. Now what is P(none defective) ? = q^144 P(1 defective) = nC1 q^143.p P(2 defective) = nC2 q^142.p^2 P(3 defective) = nC3 q^141.p^3 These are also the first four terms of the binomial expansion of (p+q)^144

OpenStudy (amistre64):

\[{144\choose 3}p^{.02}q^{.98}\] maybe?

OpenStudy (amistre64):

got my ps and qs misappropriated lol

OpenStudy (jamesj):

( I should have written 144 for n above )

OpenStudy (anonymous):

\[\dbinom{144}{3}(.02^3)(.98)^{141}\] for 3 defective, others are similar

OpenStudy (anonymous):

thats is for each one use \[\dbinom{n}{k}p^k(1-p)^{n-k}\] with \[k=0,1,2,3, p=.02,1-p=.98\]

OpenStudy (anonymous):

oh right, what jamesj said above!

OpenStudy (anonymous):

would help if i read the question.. it says "3 are defective"

OpenStudy (anonymous):

\[\dbinom{144}{3}(.02)^3(.98)^{141}\] is the one number you need

OpenStudy (anonymous):

tq! ans : 225.84